OpenStudy (anonymous):

Will give Medal! Mathis made gift bags for the students in his class for his birthday. Each gift bag contains 4 pieces of candy. If he brought 112 pieces of candy total, how many gift bags did Mathis give out? 4x = 112 108 31 28 19

OpenStudy (igreen):

Solve for 'x'.. Divide 4 to both sides..what's 112 / 4?

OpenStudy (anonymous):

28

OpenStudy (igreen):

Yep, so Mathis gave out 28 bags.

OpenStudy (anonymous):

Thank you I have some other questions. Can you answer them for me?

OpenStudy (igreen):

4 is being multiplied to 'x'..so to get 'x' by itself, we have to do the opposite. And the opposite of multiplication is division. \(4x = 112\) Divide 4 to both sides: \(\dfrac{4x}{4} = \dfrac{112}{4}\) \(\dfrac{\cancel4x}{\cancel 4} = \dfrac{112}{4}\) The 4's on the left cancel out, leaving you with: \(x = \dfrac {112}{4}\) So all you have to do is divide 112 / 4, which you came up with 28.

OpenStudy (igreen):

Sure thing, just close this one and open a new one.

OpenStudy (anonymous):

Oh ok